OpenStudy (anonymous):

Can somebody help me (puppy eyes)? Find the greatest common factor of each group of terms. 1. 48, 36x 2.42a, 28a^2 3.70x^2, 84x, 42x^3 4.24x62, 42xy^2, 66xy^3 5.60a62b^5, 140a^9, 40a^3b^6 Factor out the greatest common factor in each expression. 1.42wz+28wa 2.6x^3-12x^2+18x 3.36a^3b^6-24a^4b^2+60a^5b^3 4.(x-6)a+(x-6)b 5.(y-1)^2y+(y-1))^2z

OpenStudy (anonymous):

So many numbers my brain can't function O.o

OpenStudy (anonymous):

\(\large\textbf{Part 1}\) 1. 16 2. 14a 3. 14x ... To find the answers, one way is to factor each term completely. For example, in group 3, \[\begin{split} 70x^2 &= 2\cdot5\cdot7\cdot x\cdot x\\ 84x &= 2\cdot2\cdot3\cdot7\cdot x\\ 42x^3 &= 2\cdot3\cdot7\cdot x\cdot x\cdot x \end{split}\]You see that they have precisely one factor 2 in common, no factor 3 (because the first one doesn't have one), no factor 5, precisely one factor 7 and precisely one factor x. Multiply the common factors 2*7*x to get the answer. \(\large\textbf{Part 2}\) Do the same thing as in Part 1 except that you also divide each term by the greatest common factor. For example, in expression 1, the gcf is 14w, so the answer is \(14w(\frac{42wz}{14w}+\frac{28wa}{14w}) = 14w(3z+2a)\)

OpenStudy (anonymous):

Sorry, in 1.1 the answer should be 12, not 16.

OpenStudy (anonymous):

@JoelSjogren Thank You :*

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

@JoelSjogren I have a hard time answering number 4 and 5 on part 2 :(

OpenStudy (anonymous):

Me too sorry. The question is vague. Either (x-6)a+(x-6)b = (x-6)(a+b) is the answer or (x-6)a+(x-6)b = ax-6a+bx-6b is.

OpenStudy (ikram002p):

this is (puppy eyes) :P \(\huge\cal\color{pink}{^❤ _❤❤(✪㉨✪) _❤❤^❤_❤}\)